There are 42 houses available. Floor plan 1 is worth 100,000 and Floor plan 2 is worth 150,000. All 42 houses together are worth 6,000,000. How many of each floor plan is available?
Answer by edjones(8007)

Let x= number of floor plans selling for 100k
100x+150(42-x)=6000 Take 3 zeros off the large numbers for easier computation.
100x+6300-150x=6000
-50x=-300
x=6 for 100k
36 for 150k
.
Ed